# On the Convergence of Muon and Beyond

### v2

@@ H.1 Training on CIFAR10
The ResNet18 model He et al. (2016) undergoes pretraining on the CIFAR-10 dataset with comprehensive hyperparameter specifications provided in Table 2. For each optimizer, the learning rate is selected via a grid search over the set $\{1\times 10^{-4},5\times 10^{-4},10^{-3},5\times 10^{-3},10^{-2},5\times 10^{-2},10^{-1}\}$. To ensure a robust comparison, all experiments are repeated over five different random seeds, and we report the mean results with one standard deviation shaded.
@@
SGD Adam Muon-Type
@@ H.2 Pretraining on C4
The LLaMA2-130M modelTouvron et al. (2023) was pre-trained on the C4 (Colossal Clean Crawled Corpus) dataset Raffel et al. (2020). Identical learning rates and scheduling protocols were systematically implemented across all optimizers during the training process. Comprehensive experimental specifications are tabulated in Table 3. We use the Muon implementation from Moonlight [^1] https://github.com/MoonshotAI/Moonlight.
@@
AdamW Muon-Type

### v3

@@ H.1 Training on CIFAR10
The ResNet18 model he2016deep undergoes pretraining on the CIFAR-10 dataset with comprehensive hyperparameter specifications provided in Table 2. For each optimizer, the learning rate is selected via a grid search over the set $\{1\times 10^{-4},5\times 10^{-4},10^{-3},5\times 10^{-3},10^{-2},5\times 10^{-2},10^{-1}\}$. To ensure a robust comparison, all experiments are repeated over five different random seeds, and we report the mean results with one standard deviation shaded.
@@
SGD Adam Muon-Variant
@@ H.2 Pretraining on C4
The LLaMA2-130M modelTouvron2023Llama2O was pre-trained on the C4 (Colossal Clean Crawled Corpus) dataset raffel2020exploring. Identical learning rates and scheduling protocols were systematically implemented across all optimizers during the training process. Comprehensive experimental specifications are tabulated in Table 3. We use the Muon implementation from Moonlight [^1] https://github.com/MoonshotAI/Moonlight.
@@
AdamW Muon-Variant